Sam earned $42,000 in selling antiques. He invested part at 15% simple interest and the rest at 12%, earning $5790 per year in interest. How much did he invest at each rate?
Answer by stanbon(75887)

Sam earned $42,000 in selling antiques. He invested part at 15% simple interest and the rest at 12%, earning $5790 per year in interest. How much did he invest at each rate?
----------
EQUATIONS:
x + y = 42000
0.15x + 0.12y = 5790
----------------------
Rearrange:
15x + 12y = 579000
12x + 12y = 12*42000
-----------------------
3x = 75000
x = $25000 (amount invested at 15%)
y = 42000-25000 = $17,000 (amount invested at 12%)
===================
Cheers,
Stan H.